OpenStudy (anonymous):

Find the quotient of -36x^3-6x^2+12x over -6x −36x3 − 6x2 − 2 −6x2 + x − 2 6x2 − x + 2 6x2 + x − 2

OpenStudy (lgbasallote):

\[\frac{-36x^3 - 6x^2 +12x} {-6x} \implies \frac{-36x^3}{-6x} + \frac{-6x^2}{-6x} + \frac{12x}{-6x}\]

OpenStudy (lgbasallote):

do you get that?

OpenStudy (anonymous):

yes , thank you(:

OpenStudy (lgbasallote):

\[\huge \color{red}{\textbf{<tips hat>}}\]